Suppose that the probability of a person having a headache is 0.01, that the probability of a person having a fever given that the person has a headache is 0.4, and that the probability of a person having a fever is 0.02. Find the probability that a person has a headache given that the person has a fever.


1
Expert's answer
2021-11-10T18:37:24-0500

Let "A" denote the event "a person has a headache", "B" denote the event "a person has a fever".

Given "P(A)=0.01, P(B|A)=0.4, P(B)=0.02."

Then


"P(B|A)=\dfrac{P(A\cap B)}{P(A)}, P(A|B)=\dfrac{P(A\cap B)}{P(B)}"

"P(A|B)=\dfrac{P(B|A)P(A)}{P(B)}=\dfrac{0.4(0.01)}{0.02}=0.2"

The probability that a person has a headache given that the person has a fever is 0.2.